OpenStudy (greatlife44):

find the tangent line to the curve y = -3x^2+5x-1 (0,1)

OpenStudy (greatlife44):

Here's what I did

OpenStudy (greatlife44):

Took the derivative of the function and then evaluated it at the point x = 0 \[\frac{ dy }{ dx }(-3x^{2}+5x-1) = -6x+5\] \[|_{x=0} -6(0)+5 = 5 \] they said equation I assume i'm done here right?

OpenStudy (agent0smith):

You found the slope of the tangent line. Tangent line is an equation of a line.

OpenStudy (agent0smith):

You have the slope, and the point the tangent line passes through.

OpenStudy (greatlife44):

that's not the same as the y intercept right?

OpenStudy (agent0smith):

Use y=mx+b or y-y1 = m(x-x1). You have the slope, m. And the point the line passes through. The point happens to be the y intercept in this case.

OpenStudy (anonymous):

If \(y\) is the tangent line to the function \(f\) at \(x=a\), then: \[ \frac{y(x)-y(a)}{x-a} =f'(a) \]

OpenStudy (greatlife44):

okay, so in this case it would be y =5x+1right? is that the case in every situation?

OpenStudy (solomonzelman):

Your slope is \(\color{#000000}{ \displaystyle f'(0) }\), correct? And your points is \(\color{#000000}{ \displaystyle (0,f(0)) }\) (well, you know f(0)=1, but I will write it ge nerally) So, the equation of the tangent line is: \(\color{#000000}{ \displaystyle y-y_1=m(x-x_1)\quad \Longrightarrow \quad y-f(0)=f'(0)\cdot (x-0) }\)

OpenStudy (solomonzelman):

That is just an easy use of a point slope form.

OpenStudy (greatlife44):

yep

OpenStudy (agent0smith):

They don't always give you the y-intercept. But when they do, they prefer Dos Equis.

OpenStudy (solomonzelman):

Yes, you can certainly write the equation like you want, but if you don't want to preform any algebra in your head and had a hard day, then there you go ... ))

OpenStudy (solomonzelman):

perform**

OpenStudy (greatlife44):

interesting thanks!

OpenStudy (solomonzelman):

lol, yw.

OpenStudy (anonymous):

Eh, I my opinion the easiest thing you can do is solve for \(y\) given: \[ \frac{y-f(x_0)}{x-x_0} = f'(x_0) \]To me it's the easiest, most straight forward approach.

OpenStudy (greatlife44):

OS keeps acting up for me will look this over!